# Identification of Additive Continuous-time Systems in Open and Closed-loop

### v1

@@ 4 Additive System Identification: An Instrumental Variable Solution
This section presents the proposed instrumental variable-based method for identifying additive continuous-time systems in open or closed-loop settings.
@@
and the regressor vector $\bm{\varphi}_{i}(t_{k},\hat{\bm{\theta}}_{i})$ is expressed as
@@
Thus, by fixing $\hat{\bm{\beta}}=\hat{\bm{\beta}}^{j}$ in $\hat{\bm{\varphi}}(t_{k},\hat{\bm{\beta}})$, $\bm{\Upsilon}^{\top}(t_{k},\hat{\bm{\beta}})$ and $\bm{\varphi}^{\top}(t_{k},\hat{\bm{\beta}})$, we obtain the $j$th iteration of the proposed estimator for both open and closed-loop settings by solving for $\hat{\mathcal{B}}$ as follows:
@@
where the next iteration $\bm{\beta}^{j+1}$ is extracted from the block diagonal coefficients of $\mathcal{B}^{j+1}$ as in (30), and the instrument vector is described by (10) or (20) for the open or closed-loop setups, respectively. The iterations in (32) are computed until convergence for (8) or (11) to be satisfied.
@@
By computing the iterations in (32) with the first block of the instrument and regressor vectors given by (34) and (35) respectively, we obtain a direct extension of the proposed method for identifying marginally stable systems in additive form. This solution extends the one proposed in Solution 6.2 of [21], which is only applicable to unfactored transfer functions in closed-loop.
@@
is non-singular. Then, the converging point of the iterative procedure satisfies (8) or (11) if and only if the matrix $\bar{\mathcal{B}}$ is block-diagonal.
@@
where the last equality is due to the assumption that the limiting point $\bar{\bm{\beta}}$ satisfies the optimality condition (8) or (11) in the open or closed-loop setting, respectively. Since the modified normal matrix (36) is assumed to be non-singular, we reach $\mathbf{E}=\mathbf{0}$, which concludes the proof. $\square$
As seen in Lemma 4.4, the non-singularity of the matrix (32) is essential for the method to be well posed. This aspect, as well as the convergence of the iterations, will be studied extensively in the next section.
5 Consistency Analysis
Before presenting the main theorem on the generic consistency of the proposed method for open and closed-loop, we shall introduce some signals of interest. By substituting (4) into (24), we can express the regressor vector $\bm{\varphi}(t_{k},\bm{\beta})$ as the block vector formed by
@@ 5 Consistency Analysis
Open-loop: The input $u(t_{k})$ is persistently exciting of order no less than $2n$. Closed-loop: The input $r(t_{k})$ is persistently exciting of order no less than $2n+n_{c}$, where $n_{c}$ is the order of the discrete-time LTI controller $C_{\textnormal{d}}(q)$.
@@
Open-loop: The input $u(t_{k})$ and disturbance $v(t_{s})$ are stationary and mutually independent for all $k$ and $s$. Closed-loop: The reference $r(t_{k})$ and disturbance $v(t_{s})$ are stationary and mutually independent for all $k$ and $s$.
@@
All the zeros of the $j$th iteration of the model denominator estimate $A_{i,j}(p)$ have strictly negative real parts, with $A_{i,j}(p)$ and $B_{i,j}(p)$ being coprime. Closed-loop: Additionally, the discrete-time equivalent model estimates at each iteration are asymptotically stabilized by the controller $C_{\textnormal{d}}(q)$.
@@
Assumptions 1 to 4 are the same ones that have been considered in recent consistency analyses of open-loop and closed-loop refined instrumental variable estimators [33, 20]. We note that the stability condition in Assumption 3 is not restrictive, since the poles of the model at each iteration are typically reflected to the left half-plane if they are unstable. Cancelling the unstable poles using all-pass filters also circumvents this problem [21].
@@
    The modified normal matrix $\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}$ is generically non-singular with respect to the system and model denominator polynomials, provided that the following condition holds for each respective setting:
@@
      Open-loop:
      $\begin{split}\big{\|}&\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\Delta}^{u\top}(t_{k},\bm{\beta})\right\}\big{\|}_{2}\\ &<\sigma_{\textnormal{min}}\left(\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{u\top}(t_{k},\bm{\beta})\right\}\right),\end{split}$ (41)
@@
      Closed-loop:
      $\begin{split}\big{\|}&\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\Delta}^{r\top}(t_{k},\bm{\beta})\right\}\big{\|}_{2}\\ &<\sigma_{\textnormal{min}}\left(\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{r\top}(t_{k},\bm{\beta})\right\}\right),\end{split}$ (42)
@@
    Assume that (41) for open-loop or (42) for closed-loop is satisfied, and the iterations of the estimator converge for all $N$ sufficiently large to say, $\bar{\mathcal{B}}$, with $\bar{\mathcal{B}}$ being a block-diagonal matrix formed by $\bar{\bm{\beta}}_{i}$. Then, the true parameter vector $\bm{\beta}^{*}$ is the unique converging point of $\bar{\bm{\beta}}$ as the sample size tends to infinity.
@@
$\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}=\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})[\tilde{\bm{\varphi}}^{u}(t_{k},\bm{\beta})+\bm{\Delta}^{u}(t_{k},\bm{\beta})]^{\top}\right\}.$
Provided the condition (41) holds, due to Theorem 5.1 of [7], the perturbation matrix $\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\Delta}^{u\top}(t_{k},\bm{\beta})\right\}$ is small enough (in 2-norm) to not affect the non-singularity of the matrix $\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}$. Thus, we study the generic non-singularity of the matrix $\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{u\top}(t_{k},\bm{\beta})\right\}$ instead.
@@
where $\mathbf{S}(-B_{i}^{*},A_{i}^{*})$ and $\mathbf{S}(-B_{i},A_{i})$ are Sylvester matrices that are non-singular due to the coprimeness of the polynomials of the true system $G_{i}^{*}(p)$ and the coprimeness assumption on the model $G_{i}(p)$ [41, Lemma A3.1]. The vector $\mathbf{u}_{i}(t_{k})$ is formed by input derivatives, i.e.,
@@
$\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{u\top}(t_{k},\bm{\beta})\right\}=\bm{\mathcal{S}}\bm{\Phi}_{u}\bm{\mathcal{S}}_{*}^{\top},$ (46)
@@
$\bm{\Phi}_{u}=\mathbb{E}\left\{\begin{bmatrix}\frac{1}{A_{1}^{2}(p)}\mathbf{u}_{1}(t_{k})\\ \frac{1}{A_{2}^{2}(p)}\mathbf{u}_{2}(t_{k})\\ \vdots\\ \frac{1}{A_{K}^{2}(p)}\mathbf{u}_{K}(t_{k})\end{bmatrix}\begin{bmatrix}\frac{1}{A_{1}^{*}(p)A_{1}(p)}\mathbf{u}_{1}(t_{k})\\ \frac{1}{A_{2}^{*}(p)A_{2}(p)}\mathbf{u}_{2}(t_{k})\\ \vdots\\ \frac{1}{A_{K}^{*}(p)A_{K}(p)}\mathbf{u}_{K}(t_{k})\end{bmatrix}^{\top}\right\},$ (47)
and where $\bm{\mathcal{S}}$ and $\bm{\mathcal{S}}_{*}$ are block diagonal matrices formed by $\mathbf{S}(-B_{i},A_{i})$ and $\mathbf{S}(-B_{i}^{*},A_{i}^{*})$, respectively. Since these block Sylvester matrices are non-singular, in order for (46) to be generically non-singular, we must show that $\bm{\Phi}$ is generically non-singular. Provided the Assumptions 1, 2 and 4 hold, the generic non-singularity of such matrix follows from taking $S(q)=1$ and $\mathbf{z}_{i}(t_{k})=\mathbf{u}_{i}(t_{k})$ in Lemma A.6 in the Appendix, which concludes the proof. $\square$
@@
$\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{r\top}(t_{k},\bm{\beta})\right\}=\bm{\mathcal{S}}\bm{\Phi}_{r}\bm{\mathcal{S}}_{*}^{\top},$
where $\bm{\Phi}_{r}$ has the same form as (52) in the Appendix, but with $S(q)$ and $\mathbf{z}_{i}(t_{k})$ replaced by $S_{uo}(q)$ and $\mathbf{r}_{i}(t_{k})$, respectively. The generic non-singularity of this matrix follows from Lemma A.6.
@@
$\begin{split}\mathbb{E}&\left\{\hat{\bm{\varphi}}(t_{k},\bar{\bm{\beta}})\bm{\varphi}^{\top}(t_{k},\bar{\bm{\beta}})\right\}^{-1}\\ &\hskip 5.69046pt\times\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bar{\bm{\beta}})\left(\bm{\Upsilon}^{\top}(t_{k},\bar{\bm{\beta}})-\bm{\varphi}^{\top}(t_{k},\bar{\bm{\beta}})\bar{\mathcal{B}}\right)\right\}=\mathbf{0}.\end{split}$ (50)
Since the matrix inverse above is generically non-singular by Statement 1, we only need to analyze the second expectation. By following the derivation of the proposed estimator in (32), we find that
@@
$\mathbb{E}\left\{\hat{\bm{\varphi}}(t_{k},\bar{\bm{\beta}})\left(\sum_{i=1}^{K}[G_{i}(p,\bm{\theta}_{i}^{*})-G_{i}(p,\bar{\bm{\theta}}_{i})]u(t_{k})\right)\right\}=\mathbf{0}.$
@@ 6.1 Open-loop Experiment
where $e(t_{k})$ is a zero-mean Gaussian white noise of variance $0.02$, yielding a signal-to-noise ratio of approximately $9$ [dB]. Thirty different sample sizes are considered, ranging logarithmically from $N=2\cdot 10^{3}$ to $N=10^{5}$. For each sample size, we conduct $300$ Monte Carlo runs with varying input and noise realizations. The goal is to compare the performance of the SRIVC estimator with the proposed estimator in terms of the mean-square error of the DC gains $b_{i,0}$ related to each submodel. Both estimators initialize at a random system that satisfies $\bm{\beta}_{i}=\bm{\beta}_{i}^{*}(1+\mathcal{U}(-0.05,0.05))$, where $\mathcal{U}(a,b)$ denotes the uniform distribution with lower and upper limits $a$ and $b$, respectively. The maximum number of iterations of the algorithms is set to 100, and the relative error bound that is used as a termination rule for both iterative procedures is set to $10^{-10}$.
@@
In Figure 3 we present the MSE of each DC gain estimate as a function of the sample size. Both methods exhibit a linear decay in terms of the MSE of each parameter, which indicates that they are consistent in this scenario. This aligns with Theorem 5.5. Furthermore, we note a substantial decrease in the MSE for both $b_{2,0}$ and $b_{3,0}$ when comparing the proposed method to the SRIVC method, while remaining competitive with respect to the other parameters. The model parametrization of the proposed estimator offers the flexibility to enforce the additive system structure, resulting in more accurate estimates.

### v2

@@ 4 Additive System Identification: An Instrumental Variable Solution
This section presents the proposed instrumental variable-based method for identifying additive continuous-time systems in open or closed loop settings, and its consistency analysis.
4.1 Derivation of the instrumental variable solution
@@
and the regressor vector associated with the $i$th additive submodel, denoted by $\bm{\varphi}_{i}(t_{k},\hat{\bm{\theta}}_{i})$, is expressed as
@@
Thus, by fixing $\hat{\bm{\beta}}=\hat{\bm{\beta}}^{j}$ in $\hat{\bm{\varphi}}(t_{k},\hat{\bm{\beta}})$, $\bm{\Upsilon}^{\top}(t_{k},\hat{\bm{\beta}})$ and $\bm{\varphi}^{\top}(t_{k},\hat{\bm{\beta}})$, we obtain the $j$th iteration of the proposed estimator for both open and closed loop settings by solving for $\hat{\mathcal{B}}$ as follows:
@@
where the next iteration $\bm{\beta}^{j+1}$ is extracted from the block diagonal coefficients of $\mathcal{B}^{j+1}$ as in (27), and the instrument vector is described by (10) or (20) for the open or closed loop setups, respectively. The iterations in (29) are computed until convergence for (8) or (11) to be satisfied.
@@
By computing the iterations in (29) with the first block of the instrument and regressor vectors given by (31) and (32) respectively, we obtain a direct extension of the proposed method for identifying marginally stable systems in additive form. This solution extends the one proposed in Solution 6.2 of [24], which is only applicable to unfactored transfer functions in closed loop. We note that the user’s choice of including an integral action term in the model can be motivated by physical intuition, or by model order selection methods [31].
@@
is nonsingular. Then, the converging point of the iterative procedure satisfies (8) or (11) if and only if the matrix $\bar{\mathcal{B}}$ is block-diagonal.
@@
where the last equality is due to the assumption that the limiting point $\bar{\bm{\beta}}$ satisfies the optimality condition (8) or (11) in the open or closed-loop setting, respectively. Since the modified normal matrix (33) is assumed to be nonsingular, we reach $\mathbf{E}=\mathbf{0}$, which concludes the proof. $\square$
As seen in Lemma 4.4, the nonsingularity of the matrix (29) is essential for the method to be well posed. This aspect, as well as the convergence of the iterations, will be studied extensively in the next subsection.
4.2 Consistency Analysis
Before presenting the main theorem on the generic consistency of the proposed method for open and closed loop, we shall introduce some signals of interest. By substituting (4) into (LABEL:filteredregressorresidual), we can express the regressor vector $\bm{\varphi}(t_{k},\bm{\beta})$ as the block vector formed by
@@ 4.2 Consistency Analysis
Open loop: The input $u(t_{k})$ is persistently exciting of order no less than $2n$. Closed loop: The input $r(t_{k})$ is persistently exciting of order no less than $2n+n_{c}$, where $n_{c}$ is the order of the discrete-time LTI controller $C_{\textnormal{d}}(q)$.
@@
Open loop: The input $u(t_{k})$ and disturbance $v(t_{s})$ are quasi-stationary and mutually independent for all $k$ and $s$. Closed loop: The reference $r(t_{k})$ and disturbance $v(t_{s})$ are quasi-stationary and mutually independent for all $k$ and $s$.
@@
All the zeros of the $j$th iteration of the model denominator estimate $A_{i,j}(p)$ have strictly negative real parts, with $A_{i,j}(p)$ and $B_{i,j}(p)$ being coprime. Closed loop: Additionally, the discrete-time equivalent model estimates at each iteration are asymptotically stabilized by the controller $C_{\textnormal{d}}(q)$.
@@
Assumptions 1 to 4 are the same ones that have been considered in recent consistency analyses of open-loop and closed-loop refined instrumental variable estimators [36, 22]. Assumptions 1 and 2 can be readily satisfied if the input is white noise or a periodic signal of sufficient number of frequency lines. The stability condition in Assumption 3 is not restrictive, since the poles of the model at each iteration are typically reflected to the left half-plane if they are unstable. Cancelling the unstable poles using all-pass filters also circumvents this problem [24]. Finally, Assumption 4 avoids aliasing issues associated with the discrete-time implementation of the prefiltering steps, and is satisfied in practice if the sampling period is chosen small compared to the rise time of the system.
@@
    The modified normal matrix $\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}$ is generically nonsingular with respect to the system and model denominator polynomials, provided that the following condition holds for each respective setting:
@@
      Open loop:
      $\begin{split}\big{\|}&\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\Delta}^{u\top}(t_{k},\bm{\beta})\right\}\big{\|}_{2}\\ &<\sigma_{\textnormal{min}}\left(\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{u\top}(t_{k},\bm{\beta})\right\}\right),\end{split}$ (36)
@@
      Closed loop:
      $\begin{split}\big{\|}&\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\Delta}^{r\top}(t_{k},\bm{\beta})\right\}\big{\|}_{2}\\ &<\sigma_{\textnormal{min}}\left(\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{r\top}(t_{k},\bm{\beta})\right\}\right),\end{split}$ (37)
@@
    Assume that (36) for open loop or (37) for closed loop is satisfied, and the iterations of the estimator converge for all $N$ sufficiently large to say, $\bar{\mathcal{B}}$, with $\bar{\mathcal{B}}$ being a block-diagonal matrix formed by $\bar{\bm{\beta}}_{i}$. Then, the true parameter vector $\bm{\beta}^{*}$ is the unique converging point of $\bar{\bm{\beta}}$ as the sample size tends to infinity.
Remark 7.
The expressions in (36) and (37) give sufficient but not necessary conditions for the generic nonsingularity of $\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}$. They are satisfied in practice when the respective interpolation error $\bm{\Delta}^{u}(t_{k},\bm{\beta})$ or $\bm{\Delta}^{r}(t_{k},\bm{\beta})$ arising from constructing the filtered output derivatives is not significant, which typically occurs when the sampling period is chosen appropriately.
@@
$\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}=\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})[\tilde{\bm{\varphi}}^{u}(t_{k},\bm{\beta})+\bm{\Delta}^{u}(t_{k},\bm{\beta})]^{\top}\right\}.$
Provided the condition (36) holds, due to Theorem 5.1 of [8], the perturbation matrix $\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\Delta}^{u\top}(t_{k},\bm{\beta})\right\}$ is small enough (in 2-norm) to not affect the nonsingularity of the matrix $\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\bm{\varphi}^{\top}(t_{k},\bm{\beta})\right\}$. Thus, we study the generic nonsingularity of the matrix $\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{u\top}(t_{k},\bm{\beta})\right\}$ instead.
@@
where $\mathbf{S}(-B_{i}^{*},A_{i}^{*})$ and $\mathbf{S}(-B_{i},A_{i})$ are Sylvester matrices that are nonsingular due to the coprimeness of the polynomials of the true system $G_{i}^{*}(p)$ and the coprimeness assumption on the model $G_{i}(p)$ [44, Lemma A3.1]. The vector $\mathbf{u}_{i}(t_{k})$ is formed by input derivatives, i.e.,
@@
$\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{u\top}(t_{k},\bm{\beta})\right\}=\bm{\mathcal{S}}\bm{\Phi}_{u}\bm{\mathcal{S}}_{*}^{\top},$ (41)
@@
$\bm{\Phi}_{u}=\overline{\mathbb{E}}\left\{\begin{bmatrix}\frac{1}{A_{1}^{2}(p)}\mathbf{u}_{1}(t_{k})\\ \frac{1}{A_{2}^{2}(p)}\mathbf{u}_{2}(t_{k})\\ \vdots\\ \frac{1}{A_{K}^{2}(p)}\mathbf{u}_{K}(t_{k})\end{bmatrix}\begin{bmatrix}\frac{1}{A_{1}^{*}(p)A_{1}(p)}\mathbf{u}_{1}(t_{k})\\ \frac{1}{A_{2}^{*}(p)A_{2}(p)}\mathbf{u}_{2}(t_{k})\\ \vdots\\ \frac{1}{A_{K}^{*}(p)A_{K}(p)}\mathbf{u}_{K}(t_{k})\end{bmatrix}^{\top}\right\},$ (42)
and where $\bm{\mathcal{S}}$ and $\bm{\mathcal{S}}_{*}$ are block diagonal matrices formed by $\mathbf{S}(-B_{i},A_{i})$ and $\mathbf{S}(-B_{i}^{*},A_{i}^{*})$, respectively. Since these block Sylvester matrices are nonsingular, in order for (41) to be generically nonsingular, we must show that $\bm{\Phi}$ is generically nonsingular. Provided the Assumptions 1, 2 and 4 hold, the generic nonsingularity of such matrix follows from taking $S(q)=1$ and $\mathbf{z}_{i}(t_{k})=\mathbf{u}_{i}(t_{k})$ in Lemma A.6 in the Appendix, which concludes the proof. $\square$
@@
$\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bm{\beta})\tilde{\bm{\varphi}}^{r\top}(t_{k},\bm{\beta})\right\}=\bm{\mathcal{S}}\bm{\Phi}_{r}\bm{\mathcal{S}}_{*}^{\top},$
where $\bm{\Phi}_{r}$ has the same form as (47) in the Appendix, but with $S(q)$ and $\mathbf{z}_{i}(t_{k})$ replaced by $S_{uo}(q)$ and $\mathbf{r}_{i}(t_{k})$, respectively. The generic nonsingularity of this matrix follows from Lemma A.6.
@@
$\begin{split}\overline{\mathbb{E}}&\left\{\hat{\bm{\varphi}}(t_{k},\bar{\bm{\beta}})\bm{\varphi}^{\top}(t_{k},\bar{\bm{\beta}})\right\}^{-1}\\ &\hskip 5.69046pt\times\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bar{\bm{\beta}})\left(\bm{\Upsilon}^{\top}(t_{k},\bar{\bm{\beta}})-\bm{\varphi}^{\top}(t_{k},\bar{\bm{\beta}})\bar{\mathcal{B}}\right)\right\}=\mathbf{0}.\end{split}$ (45)
Since the matrix inverse above is generically nonsingular by Statement 1, we only need to analyze the second expectation. By following the derivation of the proposed estimator in (29), we find that
@@
$\overline{\mathbb{E}}\left\{\hat{\bm{\varphi}}(t_{k},\bar{\bm{\beta}})\left(\sum_{i=1}^{K}[G_{i}(p,\bm{\theta}_{i}^{*})-G_{i}(p,\bar{\bm{\theta}}_{i})]u(t_{k})\right)\right\}=\mathbf{0}.$
@@ 5.1 Open-loop Experiment
where $e(t_{k})$ is a zero-mean Gaussian white noise of variance $0.02$, yielding a signal-to-noise ratio of approximately $9$ [dB]. Thirty different sample sizes are considered, ranging logarithmically from $N=2\cdot 10^{3}$ to $N=10^{5}$. For each sample size, we conduct $300$ Monte Carlo runs with varying input and noise realizations. The goal is to compare the performance of the proposed estimator with the SRIVC estimator, which is the algorithm behind the tfest command in the MATLAB System Identification Toolbox [33]. To this end, we report the mean-square error of the DC gains $b_{i,0}$ related to each submodel. Both estimators initialize at a random system that satisfies $\bm{\beta}_{i}=\bm{\beta}_{i}^{*}(1+\mathcal{U}(-0.05,0.05))$, where $\mathcal{U}(a,b)$ denotes the uniform distribution with lower and upper limits $a$ and $b$, respectively. The maximum number of iterations of the algorithms is set to 100, and the relative error bound that is used as a termination rule for both iterative procedures is set to $10^{-10}$.
@@
In Figure 3 we present the MSE of each DC gain estimate as a function of the sample size. Both methods exhibit a linear decay in terms of the MSE of each parameter, which indicates that they are consistent in this scenario. This aligns with Theorem 4.5. Furthermore, we note a substantial decrease in the MSE for both $b_{2,0}$ and $b_{3,0}$ when comparing the proposed method to the SRIVC method, while remaining competitive with respect to the other parameters. The proposed estimator requires $12$ parameters to describe the system, whereas the standard SRIVC estimator considers $8$ poles and $6$ zeros, leading to a total of $15$ parameters to be estimated. The parsimony of the additive structure results in more accurate estimates.